# Efficient, simulation-free estimators of firing rates with Markovian surrogates

### v1

@@ Abstract
Spiking neural networks (SNNs) are powerful mathematical models to integrate the biological details of neural systems, but their complexity often makes them computationally expensive and analytically untractable. The firing rate of an SNN is a crucial first-order statistic for characterizing network activity. However, estimating firing rates analytically from even simplified SNN models is challenging due to 1) the intricate dependence between the nonlinear network dynamics and parameters, and 2) the singularity and irreversibility of spikes. In this Letter, we propose a class of computationally efficient, simulation-free estimators of firing rates. This is based on a hierarchy of Markovian approximations that reduces the complexity of SNN dynamics. We show that while considering firing rates alone is insufficient for accurate estimation of themselves, the information of spiking synchrony dramatically improves the estimator’s accuracy. This approach provides a practical tool for brain modelers, directly mapping biological parameters to firing rate.

### v2

@@ Abstract
Spiking neural networks (SNNs) are powerful mathematical models that integrate the biological details of neural systems, but their complexity often makes them computationally expensive and analytically untractable. The firing rate of an SNN is a crucial first-order statistic to characterize network activity. However, estimating firing rates analytically from even simplified SNN models is challenging due to 1) the intricate dependence between the nonlinear network dynamics and parameters and 2) the singularity and irreversibility of spikes. In this Letter, we propose a class of computationally efficient, simulation-free estimators of firing rates. This is based on a hierarchy of Markovian approximations that reduces the complexity of SNN dynamics. We show that while considering firing rates alone is insufficient for accurate estimations of themselves, the information of spiking synchrony dramatically improves the estimator’s accuracy. This approach provides a practical tool for brain modelers, directly mapping biological parameters to firing rate.
@@ V Discussion
VI Acknowledgment
This work was partially supported by the National Science and Technology Innovation STI2030-Major Project No. 2022ZD0204600 and the Natural Science Foundation of China through Grant No. 31771147 (Z.W., L.T.). Z.X. was supported by the Courant Institute of Mathematical Sciences at New York University during most of the time for this study, and then Shanghai Municipal Education Commission through the Eastern Talent program. We thank Professor Aaditya Rangan (New York University) for helpful discussions.